The sum of three integers in AP is 15 and their product is 80. The integers are?
Answer by stanbon(75887)

The sum of three integers in AP is 15 and their product is 80. The integers are?
-------
1st: a-d
2nd: a
3rd: a+d
--------------
Sum = 3a = 15
a = 5
-----------------
Product: a(a^2-d^2)= 80
-----------------
5(25-d^2) = 80
25-d^2 = 16
d^2 = 9
d = 3
------
1st: 5-3 = 2
2nd: 5
3rd: 5+3 = 8
----------------------
Cheers,
Stan H.
---------------------